OpenStudy (anonymous):

how to convert z = 2(cos60 + isin60) to rectangular form

OpenStudy (anonymous):

for this you only need \[\cos(60)=\frac{1}{2}\] adn \[\sin(60)=\frac{\sqrt{3}}{2}\] plug em in

OpenStudy (anonymous):

then multiply by 2 and get \[1+\sqrt{3}i\]

OpenStudy (anonymous):

going the other way is harder, but changing from trig form to rectangular means evaluate the trig functions